ServiceAgent1:

Lindsay and Lorraine are trying to match the jump rope world record. Together, they need to jump 48 times in a row. Lindsay has gotten 14 jumps in a row, and Lorraine has gotten 13. Write an equation using (j) as the variable and show your work to determine how many more jumps they need to complete.

ServiceAgent1:

@carmelle

MxxnLight:

@cArmelle

carmelle:

I think hoodmemes is doing it :)

MxxnLight:

Ah, ok-

carmelle:

ok

Hoodmemes:

Honestly as u can see it states " Lindsay has gotten 14 jumps in a row, and Lorraine has gotten 13." So u wanna (Add) (13) and (14) together \[13+14=27\] (Subtract) (27) from (48) \[48-27=21\] So they will need (21) more jumps to complete/match the record. (I feel as if something is wrong with my work,so i'd advise u to wait for a second opinion from somebody before actually trusting my answer)

Hoodmemes:

@carmelle I need your opinion on this gurl-

MxxnLight:

Hold on this slow poke

MxxnLight:

@cArmelle come back ;--;

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes Honestly as u can see it states " Lindsay has gotten 14 jumps in a row, and Lorraine has gotten 13." So u wanna (Add) (13) and (14) together \[13+14=27\] (Subtract) (27) from (48) \[48-27=21\] So they will need (21) more jumps to complete/match the record. (I feel as if something is wrong with my work,so i'd advise u to wait for a second opinion from somebody before actually trusting my answer) \(\color{#0cbb34}{\text{End of Quote}}\) That's right but, @ServiceAgent1 is it asking for the equation with the variable (ex: 1+1= J) Or do you have to solve it and then say j=21?

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes Honestly as u can see it states " Lindsay has gotten 14 jumps in a row, and Lorraine has gotten 13." So u wanna (Add) (13) and (14) together \[13+14=27\] (Subtract) (27) from (48) \[48-27=21\] So they will need (21) more jumps to complete/match the record. (I feel as if something is wrong with my work,so i'd advise u to wait for a second opinion from somebody before actually trusting my answer) \(\color{#0cbb34}{\text{End of Quote}}\) That's right but, @ServiceAgent1 is it asking for the equation with the variable (ex: 1+1= J) Or do you have to solve it and then say j=21? \(\color{#0cbb34}{\text{End of Quote}}\) like \[14 + 13 = j\]

simsharrison:

^ =27

Hoodmemes:

Thank u for checking over my work bro!

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes Thank u for checking over my work bro! \(\color{#0cbb34}{\text{End of Quote}}\) np

carmelle:

I had to leave but I think your correct, even though im 14 hours late, sorry